### v1

@@ Abstract
Context. It was the best of times, it was the worst of times, …- for the thermal/suprathermal particle populations in the vicinity of two traveling interplanetary shocks observed by Solar Orbiter on 2023-11-29 07:51:17 UTC and 2023-11-30 10:47:26 UTC at $\sim 0.83$ astronomical units from the Sun. The first (quasi-parallel) shock was an extremely interesting example of the microphysics involved in heliospheric shocks but did not accelerate a significant number of particles. The second (quasi-perpendicular) shock while not as interesting from a microphysics point of view, did accelerate particles - likely at a magnetically connected remote location.
Aims. We investigate these two very dissimilar shocks and elucidate their non-equilibrium features. We do not provide explanations of all observed features, our aim is to report them here for future reference.
Methods. We use high-resolution data obtained with Solar Orbiter’s Energetic Particle Detector (EPD), magnetometer (MAG), and Solar Wind Analyzer (SWA) to exhibit the very different natures of these two shocks and describe the detailed properties of suprathermal and energetic particles in their vicinity.
Results. We observe very different behavior of the energetic particle population because the two shocks are quite different. Solar wind protons and $\alpha$-particles are highly dynamic at the first, their beams appear to align well with rapid oscillations of the interplanetary magnetic field. Suprathermal particles associated with the second shock exhibit clear non-equilibrium and anisotropic features in their differential intensities at time scales comparable to the proton gyroperiod.
Conclusions. The different geometries of the two shocks resulted in highly dissimilar populations of suprathermal and energetic particles in their vicinity. The first shock was associated with very interesting microphysics of the bulk plasma velocity distribution, the second resulted in similarly interesting microphysics of the accelerated particles. Both showed strong temporal variability of the particle populations at scales comparable to the proton gyroperiod.
@@ 1 Introduction
The differential flux of energetic particles in the vicinity of traveling interplanetary (IP) shocks is believed to be determined by a number of factors (see e.g., Desai & Giacalone, 2016, for a review). Energetic particles are thought to be accelerated out of a pool of suprathermal particles; physical and geometrical properties of the shock (such as compression ratio and angle of the shock normal) are expected to play an important role in the actual acceleration of the particles (see, e.g., Kartavykh et al., 2025, for a very recent review). Here we report on two shocks that encountered a very dissimilar pool of suprathermal particles and were separated in time by a little under 27 hours when Solar Orbiter was at $\sim 0.83$ astronomical units from the Sun. Their properties are reported in tab. 1 and discussed in sec. 2. The first (quasi-parallel shock) did not appear to accelerate a significant number of particles, while the second (quasi-perpendicular shock) did have a noticeable effect. Observations around the first shock show highly correlated behavior of waves and bulk plasma at the time scale of seconds. Observations around the second shock at the same high time resolution show extremely rich fine st ructure and non-equilibrium distributions of energetic particles.
@@ 2 Overview of the time period
Figure 1 presents an overview of the observations and shows, from top to bottom, the magnitude of the interplanetary magnetic field (IMF, in black) and radial solar wind speed ($V_{R}$, in blue) solar wind proton density (black) and temperature (red, in $kT$ energy equivalent), color-coded differential intensities as measured by the sun, anti-sun, north, and south telescopes of EPT. The central pitch angles of the four telescopes are overplotted with the angle information shown on the right-hand axis. The bottom two panels show the IMF azimuth (B_az) and elevation (B_ele) angle (both in black), as well as the tangential and normal components of solar wind speed ($V_{T}$ and $V_{N}$, both in blue). The two shocks are clearly visible in the IMF magnitude, they are indicated by vertical red lines at 2023-11-29 07:51:17 and 2023-11-30 10:47:26. All times are given in UTC in this manuscript. Key parameters of the two shocks are presented in tab. 1. Here, the shock normal has been computed using the Mixed Mode 3 method (MX3 Paschmann & Schwartz, 2000), and the shock speed using mass flux conservation. An ensemble of upstream/downstream averaging windows ranging from 10 seconds to 5 minutes has been used to evaluate the shock parameters (see Trotta et al., 2022, for a full description). About two hours before the first shock (around 06:00), EPT’s sun and south telescopes begin to see elevated particle fluxes at energies below 200 keV. Before that time, that low-energy pool population appears to be isotropic and less populated. Because the sun and south telescopes cover a similar range of pitch angles around 06:30 on 2023-11-29 they both observe the same elevated particle flux. The particle population is strongly anisotropic downstream of the first shock and streams away from the Sun. It is strongly reminiscent of a hemispheric distribution as witnessed by the oscillating intensities in the north and south telescopes that coincide with the times when the central pitch angles of these telescopes lie around 90 degrees. Particles appear to be more isotropically distributed at the shock ramp and downstream of the second shock. The second shock (2023-11-30 10:47:26) is stronger than the first and obviously has a much stronger influence on the energetic particle population which appears to be more isotropic upstream and downstream of shock #2 than at shock #1.
@@ 3 Shock # 1
Properties of the IMF surrounding shock #1 are investigated in more detail in fig. 2. The shock is preceded by a region of low variability and a systematic rotation of the IMF between  07:00 and the time of the shock which can easily be seen in the top panel. This could indicate that Solar Orbiter was immersed in a small interplanetary coronal mass ejection (ICME) at this time, but its identification is not the topic of this paper. Nevertheless, this would explain the low level of IMF fluctuations which are also visible in the 2nd panel from the top. Analyzing the trace wavelet spectrogram in the mid-panel of fig. 2, low levels of magnetic fluctuations are observed during this time period. Closer to the shock (indicated by a vertical solid red line), starting at 7:50:00, enhanced wave activity can be seen at high frequencies ($\gtrsim 1$ Hz). The bottom panel shows the normalized magnetic helicity, defined as $\sigma_{m}=\frac{2\Im\left(\tilde{B_{T}^{\star}}\tilde{B_{N}}\right)}{|\tilde{B_{R}}|^{2}+|\tilde{B_{T}}|^{2}+|\tilde{B_{N}}|^{2}}$, where $\mathrm{B}$ indicates the magnetic field components, the $\tilde{}$ represents the wavelet- transformed quantities and $\star$ represents the complex conjugation operation (Matthaeus et al., 1982). Upstream of the shock, we observe a clear signature of consistently high $\sigma_{m}$ at ion scales, often observed in the inner heliosphere (Verniero et al., 2020). These waves are visible as the red patches at the ion plasma frequency which is shown as the black line in the bottom panel of fig. 2. Note that these frequencies are not Doppler shifted, and a detailed investigation of such waves and their effect on the suprathermal particles upstream of shock #1 will be part of a separate study. Here, we note that this wave activity disappears downstream of the shock, a behavior similar to what recently observed upstream of an inner heliospheric shock (Trotta et al., 2024) and, interestingly, in a weak shock propagating in CME material as observed by Trotta et al. (2024)b. In that scenario of CME-CME interaction, compatible with the one analyzed here, the low upstream level of fluctuations allows waves and shock-reflected particles to travel very far upstream of the shocks without undergoing significant scattering. This very likely explains the long-lasting ($\sim$ 1 hour) waves upstream of the first shock and the streaming particles discussed in sec. 2.
@@
Data from the Proton-Alpha Sensor (PAS) of the SWA instrument were acquired in an experimental mode to obtain reduced velocity distribution functions (VDFs) at a cadence of one second instead of the usual four seconds. This was achieved by reducing the number of elevation angles (and energy-per-charge steps) from 9 (96) to 5 (48). To ensure that the proton peak is not lost, a full scan was performed every 100 seconds. This new operation mode can result in somewhat spiky proton moments. We compared full $E/q$-spectra at 1-s and 100-s and found that the protons were accurately tracked most of the time during the time period of interest, and that the proton VDFs were completely captured. Solar wind properties from MAG and PAS are shown in fig. 4 with magnetic field data in panel a), an ion omnidirectional energy flux spectrogram in panel b), proton density and temperature in panel c), proton radial velocity and total velocity in panel d), and proton tangential (T) and normal (N) velocities in panel e). The spectrogram shows that PAS captured the proton VDF well, despite the restricted (but higher cadence) data product. Vertical dashed lines indicate the times for which VDFs are shown in fig. 5. Interestingly, around the shock time between 07:51:00 and 07:51:15 in fig. 4, the proton tangential and normal velocities show fluctuations similar to those of the magnetic fields, which results in the fluctuations of the proton moments seen in panel c).
@@
Figure 6 shows the same time period as fig. 4 with shock #1 again indicated by a vertical red line. The format is the same as in fig. 1 with the omission of solar wind density and temperature and the addition of a panel that shows the pitch-angle distribution of $\sim$56 keV protons as observed by the four apertures of EPT. The already discussed strong wave activity is seen upstream of the shock, but the short-period ion-cyclotron waves are out of resonance with energetic and suprathermal protons. Low-energy protons are seen in the sun-telescope before and after the shock without any clear discontinuity in their flux. The third panel from the bottom shows color-coded differential intensity in the EPD energy bin $53.8-58.7$ keV. The flux is concentrated at pitch angles $<90^{\circ}$, indicating a strongly anisotropic pitch-angle distribution. Thus, despite the very interesting kinetic physics (affecting the thermal (bulk and beam) solar wind) visible at shock #1, it apparently did not strongly affect energetic protons (i.e., with energies greater than $\sim 25$ keV). Nevertheless, the flux of energetic protons is highly anisotropic. The flux of electrons in this event is too low to be measurable by EPT at these time scales and is therefore not shown.
@@ 4.2 Time period 1)
Figure 8 zooms in to time period 1 in a similar format as fig. 1 but shows only the sunward telescope of EPT (3rd panel from the top) and, in addition, data from the central $3\times 3$ pixels of STEP (4th panel from the top). A ”loop”-like feature is clearly seen between 10:38 and just beyond 10:39 between the second and third vertical blue lines (counted from the left). It clearly extends from STEP energies to EPT energies but is in fact extremely anisotropic, as can be seen in fig. 13 in the appendix. The velocity distribution between the second and third vertical dashed blue lines is obviously unstable, but lasts only for about one minute. It ends at the magnetic discontinuity (3rd blue line), but extends beyond that at higher energies. Nevertheless, it persists for several proton gyro-periods ($\tau_{p}\sim 10$ seconds), thus providing a limit on the growth rate of instabilities which drive the system towards equilibrium. It is worth noting that such features would not have been recognized with traditional particle instruments because of their much lower time and energy resolution.
@@ 4.4 Time period 3)
Figure 10 is in the usual format, but differential intensities have been multiplied by energy to more clearly show the many substructures seen in time periods 3) - 5). The shock is indicated by the vertical red line, the vertical dashed blue lines are discussed in subsection 4.6. The feature at time period 3) can be seen in the northward-pointing telescope between 10:46:40 and 10:46:57 and appears to show velocity dispersion. It begins at about the time where fluctuations in the IMF begin to be visible and probably only appears to end because the energy of this ”beam” was too low to be measured by EPT. As STEP does not have a telescope in the north direction, we could not measure this putative extension. This dispersive feature is seen at small pitch angles, i.e., points along the IMF and precedes the shock by approximately one minute. We interpret this as a consequence of irregular injection at a remote location along the shock surface, similar to what has been reported by Trotta et al. (2023) and Yang et al. (2023, 2024), but more pronounced. Additional dispersive features are seen at similarly low energies in the anti-sun telescope immediately ahead of the shock during the prominent IMF fluctuations (even when the pitch angle scanned by this telescope is around 90 deg).
@@ 5 Discussion
The two events analyzed here provide an excellent opportunity to study a variety of processes of fundamental importance for energy conversion in space and astrophysical plasmas. In particular, shock #1 shows a plethora of small-scale features, namely shock–induced upstream fluctuations and strongly out-of-equilibrium velocity distribution functions that underline the importance of ion kinetics in reconstructing how energy is converted at collisionless shocks. Such processes, widely addressed at Earth’s bow shock (e.g., Schwartz et al., 2022), remain elusive at IP shocks due to technical challenges, for which Solar Orbiter is advancing the state-of-the-art (see Dimmock et al., 2023, for example). Further, as efficient proton acceleration is observed at shock #2, this study elucidates how preceding events could play a role in pre-conditioning plasma for further acceleration. Such multiple events are more common and may be important in conditions of solar activity maximum (see Trotta et al., 2025, for an overview). However, to pin-point the role of preceding events in the generation of particle distributions that can be easily re-accelerated to higher energies, input from numerical modeling would be crucial (see Nyberg et al., 2024).

### v2

@@ Abstract
Context. Energetic particles in interplanetary space are normally measured at time scales that are long compared to the ion gyroperiod. Such observations by necessity average out the microphysics associated with the acceleration and transport of 10s - 100s keV particles.
Aims. We investigate previously unseen non-equilibrium features that only become observable at very high time resolution, and discuss possible explanations of these features.
Methods. We use unprecedentedly high-time-resolution data that were acquired by the in situ instruments on Solar Orbiter in the vicinity of two interplanetary shocks observed on 2023-11-29 07:51:17 UTC and 2023-11-30 10:47:26 UTC at $\sim 0.83$ astronomical units from the Sun.
Results. The solar-wind proton beam population follows the magnetic field instantaneously, on time scales which are significantly shorter than a gyro-period. Energetic particles, despite sampling large volumes of space, vary on remarkably short time scales, typically on the order of the convection time of their gyro-radius. Non-equilibrium features such as bump-on-tail distributions of energetic particles are formed by small-scale magnetic structures in the IMF.
Conclusions. High-time-resolution observations show previously unobserved microphysics in the vicinity of two traveling interplanetary shocks, including ion reflection at a current sheet, which may explain where ions are reflected in shock acceleration.
@@ 1 Introduction
The differential flux of energetic particles in the vicinity of traveling interplanetary (IP) shocks is believed to be determined by a number of factors (see e.g., Desai & Giacalone, 2016, for a review). Particles are accelerated by two main mechanisms depending on the orientation of the interplanetary magnetic field (IMF) with respect to the shock normal. It has thus become customary to distinguish between (quasi-) parallel and (quasi-) perpendicular shocks. In the (quasi-) parallel case, waves upstream and downstream of the shock scatter particles back and forth across the shock and thus accelerate particles through the first-order Fermi process Fermi (1949). The detailed theory of this process using a diffusive approach was worked out by, e.g., Jokipii (1966); Fisk (1971); Axford et al. (1977). Early observations at quasi-perpendicular shocks showed that particles gain energy by (gradient) drifting along the induced electric field at the shock surface (Sarris & van Allen, 1974; Pesses et al., 1979; Tsurutani et al., 1982, see, e.g.,), the theory was developed by Hudson (1965). Both processes can be treated as a diffusive process and included in the particle transport equation Parker (1965); Desai & Giacalone (2016). Energetic particles are thought to be accelerated out of a pool of suprathermal particles; physical and geometrical properties of the shock (such as compression ratio and angle of the shock normal) are expected to play an important role in the actual acceleration of the particles. A large body of literature exists about particles accelerated at interplanetary shocks, ranging from detailed descriptions of one single shock (e.g., the November 12, 1978 quasi-parallel shock Kennel et al., 1984b, a; Tsurutani et al., 2024) to comprehensive statistical studies of shock properties (see, e.g., Trotta et al., 2025, for a very recent review of shocks encountered by Solar Orbiter). Kennel et al. (1985) and Papadopoulos (1985) give two excellent tutorials of the physics of particle acceleration at shocks, of which Kartavykh et al. (2025) provide a very recent review.
Here we report on two shocks that encountered a very dissimilar pool of suprathermal particles and were separated in time by a little under 27 hours when Solar Orbiter was at $\sim 0.83$ astronomical units from the Sun. Their properties are reported in tab. 1 and discussed in sec. 2. As will be shown in sec. 3 and 4 the first (quasi-parallel shock) did not appear to accelerate a significant number of particles, while the second (quasi-perpendicular shock) did have a noticeable effect. Observations around the first shock show highly correlated behavior of waves and bulk plasma at the time scale of seconds. Observations around the second shock at the same high time resolution show extremely rich fine structure and non-equilibrium distributions of energetic particles.
@@ 2 Overview of the time period
Figure 1 presents an overview of the observations and shows, from top to bottom, the magnitude of the interplanetary magnetic field (IMF, in black) and radial solar wind speed ($V_{R}$, in blue) solar wind proton density (black) and temperature (red, in $kT$ energy equivalent), color-coded differential intensities as measured by the sun, anti-sun, north, and south telescopes of EPT. The central pitch angles of the four telescopes are overplotted with the angle information shown on the right-hand axis. The bottom two panels show the IMF azimuth (B_az) and elevation (B_ele) angle (both in black), as well as the tangential and normal components of solar wind speed ($V_{T}$ and $V_{N}$, both in blue). The two shocks are clearly visible in the IMF magnitude, they are indicated by vertical red lines at 2023-11-29 07:51:17 and 2023-11-30 10:47:26. All times are given in UTC in this manuscript. Key parameters of the two shocks are presented in tab. 1. They differ slightly from the values given in the SERPENTINE data center https://doi.org/10.5281/zenodo.12518015 because here we used averaging intervals which were tailored to these two specific shocks whereas the values in the SERPENTINE data base were derived using similar windows for all shock. The shock normal has been computed using the Mixed Mode 3 method (MX3 Paschmann & Schwartz, 2000), and the shock speed using mass flux conservation. An ensemble of upstream/downstream averaging windows ranging from 10 seconds to 5 minutes has been used to evaluate the shock parameters (see Trotta et al., 2022, for a full description). Both shocks are supercritical as will become obvious when we present them in more detail in the following sections 3 and 4. Both shocks show the ”foot” and overshoots which are typical of supercritical shocks Bavassano-Cattaneo et al. (1986).
About two hours before the first shock (around 06:00), EPT’s sun and south telescopes begin to see elevated particle fluxes at energies below 200 keV. Before that time, that low-energy pool population appears to be isotropic and less populated. Because the sun and south telescopes cover a similar range of pitch angles around 06:30 on 2023-11-29 they both observe the same elevated particle flux. The particle population is strongly anisotropic downstream of the first shock and streams away from the Sun. It is strongly reminiscent of a hemispheric distribution as witnessed by the oscillating intensities in the north and south telescopes that coincide with the times when the central pitch angles of these telescopes lie around 90 degrees. Particles appear to be more isotropically distributed at the shock ramp and downstream of the second shock. The second shock (2023-11-30 10:47:26) is stronger than the first and obviously has a much stronger influence on the energetic particle population which appears to be more isotropic upstream and downstream of shock #2 than at shock #1.
@@ 3 Shock # 1
Properties of the IMF surrounding shock #1 are investigated in more detail in fig. 2. The shock is preceded by a region of low variability and a systematic rotation of the IMF between  07:00 and the time of the shock which can easily be seen in the top panel. This indicates that Solar Orbiter was immersed in a small interplanetary coronal mass ejection (ICME) at this time, but its identification is not the topic of this paper. Nevertheless, this would explain the low level of IMF fluctuations which are also visible in the 2nd panel from the top. Analyzing the trace wavelet spectrogram in the mid-panel of fig. 2, low levels of magnetic fluctuations are observed during this time period. Closer to the shock (indicated by a vertical solid red line), starting at 7:50:00, enhanced wave activity can be seen at high frequencies ($\gtrsim 1$ Hz). The bottom panel shows the normalized magnetic helicity, defined as $\sigma_{m}=\frac{2\Im\left(\tilde{B_{T}^{\star}}\tilde{B_{N}}\right)}{|\tilde{B_{R}}|^{2}+|\tilde{B_{T}}|^{2}+|\tilde{B_{N}}|^{2}}$, where $\mathrm{B}$ indicates the magnetic field components, the $\tilde{}$ represents the wavelet- transformed quantities and $\star$ represents the complex conjugation operation (Matthaeus et al., 1982). Upstream of the shock, we observe a clear signature of consistently high $\sigma_{m}$ at ion scales, often observed in the inner heliosphere (Verniero et al., 2020). These waves are visible as the red patches at the ion plasma frequency which is shown as the black line in the bottom panel of fig. 2. Note that these frequencies are not Doppler shifted, and a detailed investigation of such waves and their effect on the suprathermal particles upstream of shock #1 will be part of a separate study. Here, we note that this wave activity disappears downstream of the shock, a behavior similar to what recently observed upstream of an inner heliospheric shock (Trotta et al., 2024) and, interestingly, in a weak shock propagating in CME material as observed by Trotta et al. (2024)b. In that scenario of CME-CME interaction, compatible with the one analyzed here, the low upstream level of fluctuations allows waves and shock-reflected particles to travel very far upstream of the shocks without undergoing significant scattering. This very likely explains the long-lasting ($\sim$ 1 hour) waves upstream of the first shock and the streaming particles discussed in sec. 2.
@@
Data from the Proton-Alpha Sensor (PAS) of the SWA instrument were acquired in an experimental mode to obtain reduced velocity distribution functions (VDFs) at a cadence of one second instead of the usual four seconds. This was achieved by reducing the number of elevation angles (and energy-per-charge steps) from 9 (96) to 5 (48). To ensure that the proton peak is not lost, a full scan was performed every 100 seconds. This new operation mode can result in somewhat spiky proton moments. We compared full $E/q$-spectra at 1-s and 100-s and found that the protons were accurately tracked most of the time during the time period of interest, and that the proton VDFs were completely captured. Solar wind properties from MAG and PAS are shown in fig. 4 with magnetic field data in panel a), an ion omnidirectional energy flux spectrogram in panel b), proton density and temperature in panel c), proton radial velocity and total velocity in panel d), and proton tangential (T) and normal (N) velocities in panel e). The spectrogram shows that PAS captured the proton VDF well, despite the restricted (but higher cadence) data product. Vertical dashed lines indicate the times for which VDFs are shown in fig. 5. In panel a) one can clearly see the waves emanating ahead of the shock which indicate that the shock is supercritical Bavassano-Cattaneo et al. (1986), as we mentioned in sec. 2. Moreover, as can be seen from the VDFs collected at line 5, strong ion reflection is also found, a “smoking gun” of the shocks’ supercriticality and rarely resolved at IP shocks because of instrumental limitations (see Dimmock et al., 2023). Interestingly, around the shock time between 07:51:00 and 07:51:15 in fig. 4, the proton tangential and normal velocities show fluctuations similar to those of the magnetic fields, which results in the fluctuations of the proton moments seen in panel c).
@@
Figure 6 shows the same time period as fig. 4 with shock #1 again indicated by a vertical red line. The format is the same as in fig. 1 with the omission of solar wind density and temperature and the addition of a panel that shows the pitch-angle distribution of $\sim$56 keV protons as observed by the four apertures of EPT. The already discussed strong wave activity is seen upstream of the shock, but the short-period ion-cyclotron waves are out of resonance with energetic and suprathermal protons. Low-energy protons are seen in the sun-telescope before and after the shock without any clear discontinuity in their flux. The third panel from the bottom shows color-coded differential intensity in the EPD energy bin $53.8-58.7$ keV. The flux is concentrated at pitch angles $<90^{\circ}$, indicating a strongly anisotropic pitch-angle distribution. Thus, despite the very interesting kinetic physics (affecting the thermal (bulk and beam) solar wind) visible at shock #1, it apparently did not strongly affect energetic protons (i.e., with energies greater than $\sim 25$ keV). Nevertheless, the flux of energetic protons is highly anisotropic. The flux of electrons in this event is too low to be measurable by EPT at these time scales and is therefore not shown. Based on the properties of this shock (see tab. 1), we would expect it to accelerate both electrons and ions. Tsurutani & Lin (1985) found that a minimum shock speed along the upstream IMF of 250 km/s is needed for a shock to accelerate $\geq 2$ keV electrons and $\geq 47$ keV ions. Shock #1 clearly exceeds this limit.
As discussed above, shock #1 is preceded by a small CME and its low level of fluctuations in the IMF. This likely also contributes to the low acceleration efficiency of this shock because the low level of upstream turbulence (inside the CME) may not be sufficient to scatter particles back towards the shock.
@@ 4.2 Time period 1)
Figure 8 zooms in to time period 1 in a similar format as fig. 1 but shows only the sunward telescope of EPT (3rd panel from the top) and, in addition, data from the central $3\times 3$ pixels of STEP (4th panel from the top). A ”loop”-like feature is clearly seen between 10:38 and just beyond 10:39 between the second and third vertical blue lines (counted from the left). It clearly extends from STEP energies to EPT energies but is in fact extremely anisotropic, as can be seen in fig. 13 in appendix B. The velocity distribution between the second and third vertical dashed blue lines is obviously unstable, but lasts only for about one minute. It ends at the magnetic discontinuity (3rd blue line), but extends beyond that at higher energies. Nevertheless, it persists for several proton gyro-periods ($\tau_{p}\sim 10$ seconds), thus providing a limit on the growth rate of instabilities which drive the system towards equilibrium. It is worth noting that such features would not have been recognized with traditional particle instruments because of their much lower time and energy resolution.
@@ 4.4 Time period 3)
Figure 10 is in the usual format, but differential intensities have been multiplied by energy to more clearly show the many substructures seen in time periods 3) - 5). The shock is indicated by the vertical red line, the vertical dashed blue lines are discussed in subsection 4.6. The overshoot features typical of supercritical shocks Bavassano-Cattaneo et al. (1986) are also present but not as obvious as for shock #1 because of the more compressed $x$ axis, the larger $\theta_{Bn}$ and the higher shock speed in the spacecraft rest frame (see Table 1). The feature at time period 3) can be seen in the northward-pointing telescope between 10:46:40 and 10:46:57 and appears to show velocity dispersion. It begins at about the time where fluctuations in the IMF begin to be visible and probably only appears to end because the energy of this ”beam” was too low to be measured by EPT. As STEP does not have a telescope in the north direction, we could not measure this putative extension. This dispersive feature is seen at small pitch angles, i.e., points along the IMF and precedes the shock by approximately one minute. We interpret this as a consequence of irregular injection at a remote location along the shock surface, similar to what has been reported by Trotta et al. (2023) and Yang et al. (2023, 2024), but more pronounced. Additional dispersive features are seen at similarly low energies in the anti-sun telescope immediately ahead of the shock during the prominent IMF fluctuations (even when the pitch angle scanned by this telescope is around 90 deg).
@@ 5 Discussion
While the study of collisionless shocks in the heliosphere is a mature field (Kennel et al., 1985; Papadopoulos, 1985, e.g.,), new measurements at unprecedented time and energy resolution in the inner heliosphere are providing new and exciting insights into the microphysics of the processes associated with particle acceleration, so far mostly tested at the Earth’s bow shock only. The two events analyzed here provide an excellent opportunity to study a variety of processes of fundamental importance for energy conversion in space and astrophysical plasmas. In particular, shock #1 shows a plethora of small-scale features, namely shock–induced upstream fluctuations and strongly out-of-equilibrium velocity distribution functions that underline the importance of ion kinetics in reconstructing how energy is converted at collisionless shocks. Such processes, widely addressed at Earth’s bow shock (e.g., Schwartz et al., 2022), remain elusive at IP shocks due to technical challenges, for which Solar Orbiter is advancing the state-of-the-art (see Dimmock et al., 2023, for example). Further, as efficient proton acceleration is observed at shock #2, this study elucidates how preceding events could play a role in pre-conditioning plasma for further acceleration. Such multiple events are more common and may be important in conditions of solar activity maximum (see Trotta et al., 2025, for an overview). However, to pin-point the role of preceding events in the generation of particle distributions that can be easily re-accelerated to higher energies, input from numerical modeling would be crucial (see Nyberg et al., 2024). The contrasting observation that shock #1 had virtually no effect on energetic particles underlines this point because it was preceded by a small CME with its low level of fluctuations. This compounds the problem for particles to cross the 90 pitch angle to be reflected back into the shock for further acceleration. As was pointed out by Tsurutani et al. (2024), it is well known that there is no resonance with particles at 90 in the energy range considered in this study Horne et al. (2003). The low level of fluctuations inside the CME further enhances this problem. Thus, this shock (#1) is an extreme, limiting example of a shock that does not accelerate particles.